# Revisiting PSF models: unifying framework and high-performance implementation

### v1

@@ 3.3 Vectorial Models
As the electric field is a vectorial quantity, vectorial propagation models are necessary to accurately account for the propagation and crosstalk between the different components of the vector field. Employing these precise vectorial models is crucial for high-NA systems in which case the need to consider high angles arises.
In the vectorial model, the far field $\mathbf{e}_{\infty}(\mathbf{s})$ now has a more complex dependence on the incident field $\mathbf{e}_{\mathrm{inc}}(\mathbf{s})$ as we need to perform the basis change from a cylindrical to a spherical coordinate system:
@@
$+\left[\begin{array}[]{c}(-q_{s}+q_{p}\cos\theta)\sin 2\phi\\ q_{s}(1+\cos 2\phi)+q_{p}(1-\cos 2\phi)\cos\theta\\ -2q_{p}\sin\phi\sin\theta\end{array}\right]\frac{e_{\mathrm{inc}}^{y}(\theta,\phi)}{2}$ (13)
where $\mathbf{e}_{\mathrm{inc}}=[e_{\mathrm{inc}}^{x},e_{\mathrm{inc}}^{y},0]$. Fresnel transmission coefficients $q_{s}$ and $q_{p}$ have been introduced to account for partial reflection at interfaces, which depend on the polarization state and incidence angle. For each of the $s$ and $p$ polarizations, they correspond to the product of all transmission coefficients for each interface from medium $j$ to $j+1$:
$q_{s}^{j}=\frac{2n_{j}\cos{\theta_{j}}}{n_{j}\cos\theta_{j}+n_{j+1}\cos{\theta_{j+1}}},$ (14)
$q_{p}^{j}=\frac{2n_{j}\cos{\theta_{j}}}{n_{j+1}\cos\theta_{j}+n_{j}\cos{\theta_{j+1}}}.$ (15)
The Cartesian parametrization of the vectorial model consists of the following integral:
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{s_{x}^{2}+s_{y}^{2}\leq s_{\max}^{2}}\frac{\mathbf{e}_{\infty}(s_{x},s_{y})}{s_{z}}\exp{\left\{\mathrm{i}ks_{z}z\right\}}\exp{\left\{\mathrm{i}k(s_{x}x+s_{y}y)\right\}}\mathrm{d}s_{x}\mathrm{d}s_{y},$ (16)
which essentially boils down to computing the inverse Fourier Transform of $(\mathbf{e}_{\infty}(s_{x},s_{y})\exp{\left\{\mathrm{i}ks_{z}z\right\}}/s_{z})$, similar to the scalar case.
Using coordinate transformations similar to the scalar case, we can derive the spherical parametrization of the field in the focal region:
@@
$\exp{\left\{\mathrm{i}k\rho\sin\theta\cos(\phi-\varphi)\right\}}\exp{\left\{\mathrm{i}kz\cos\theta\right\}}\sin\theta.$ (17)
Inserting (9) into (3.3) and using the axisymmetric assumption of the incident field, we can obtain a simplified expression for the focal field as follows:
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2}\left[\begin{array}[]{c}[I_{0}^{x}-I_{2}^{x}\cos 2\varphi]-I_{2}^{y}\sin 2\varphi\\ -I_{2}^{x}\sin 2\varphi+[I_{0}^{y}+I_{2}^{y}\cos 2\varphi]\\ -2\mathrm{i}I_{1}^{x}\cos\varphi-2\mathrm{i}I_{1}^{y}\sin\varphi\end{array}\right]$ (18)
@@
$I_{0}^{a}(\rho,z)=\int_{0}^{\theta_{\max}}e_{\mathrm{inc}}^{a}(\theta)\sin\theta(\cos\theta+1)J_{0}(k\rho\sin\theta)\exp{\left\{\mathrm{i}kz\cos\theta\right\}}\mathrm{d}\theta$ (19)
$I_{1}^{a}(\rho,z)=\int_{0}^{\theta_{\max}}e_{\mathrm{inc}}^{a}(\theta)\sin^{2}\theta J_{1}(k\rho\sin\theta)\exp{\left\{\mathrm{i}kz\cos\theta\right\}}\mathrm{d}\theta$ (20)
$I_{2}^{a}(\rho,z)=\int_{0}^{\theta_{\max}}e_{\mathrm{inc}}^{a}(\theta)\sin\theta(\cos\theta-1)J_{2}(k\rho\sin\theta)\exp{\left\{\mathrm{i}kz\cos\theta\right\}}\mathrm{d}\theta$ (21)
with $a\in\{x,y\}$ and $\mathbf{e}_{\mathrm{inc}}(\theta)=[e_{\mathrm{inc}}^{x}(\theta),e_{\mathrm{inc}}^{y}(\theta),0]$.
@@ 3.4 Correction Factors
Precise PSF models commonly include several physical effects which may affect the PSF. They are added as amplitude factors $a(\mathbf{s})$ and phase factors $W(\mathbf{s})$ in the original integral over solid angles in Eq. 1:
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{\Omega}a(\mathbf{s})\exp{\left\{\mathrm{i}W(\mathbf{s})\right\}}\mathbf{e}_{\infty}(\mathbf{s})\exp{\left\{\mathrm{i}k\mathbf{s}\cdot\boldsymbol{\rho}\right\}}\mathrm{d}\Omega.$ (22)
Eq. (22) enables us to express these correction factors with full generality: in both vectorial and scalar models, for both Cartesian and spherical parametrizations. We present a detailed list of these correction factors in Section 3.4.1-3.4.4 and a graphical description of them in Figure 2.
Figure 2: Correction factors and associated physical effects. Top: phase correction factors, either introduced on the incident field or due to refraction at different planes. Bottom: amplitude correction factors, to model an incident envelope or apodization factor for energy conservation.
3.4.1 Aberrations Due to Refractive Index Mismatch
Microscopes typically have stratified layers of different refractive indices. The biological sample is usually aqueous, on top of which we place a coverslip made of glass, and the whole sample is then put in a water or oil immersion medium to increase the numerical aperture. The microscope objectives are designed to provide aberration-free images in a specific setting with design values for refractive indices and thicknesses of the different layers. Any mismatch introduces spherical aberrations due to refraction at the different layers. These aberrations can be computed using the following formula [^3] Correction factors will be expressed in spherical coordinates. $\sin\theta$ is computed in the Cartesian case using $\sin\theta=\frac{NA}{n_{i}}\sqrt{s_{x}^{2}+s_{y}^{2}}$.:
@@ 3.4.1 Aberrations Due to Refractive Index Mismatch
$\quad\left.+t_{g}\sqrt{n_{g}^{2}-n_{i}^{2}\sin^{2}\theta}-t_{g}^{*}\sqrt{\left.n_{g}^{*}\right.^{2}-n_{i}^{2}\sin^{2}\theta}\right)$ (23)
where $n_{s}$, $n_{i}$, $n_{g}$ are the refractive indices of the sample, immersion medium, and glass respectively, $t_{s}$, $t_{i}$, $t_{g}$ are the thicknesses of the sample, immersion medium, and glass respectively, and their counterparts with stars are the design conditions.
A particular case proposed in [19] is commonly encountered: it is difficult in practice to assess the thickness of the immersion medium. Since this distance is manually tuned to obtain an optimal focus of a point emitter at depth $t_{s}$ on the camera, this focusing condition gives the following relation:
@@
which can be inserted in Eq. 3.4.1. This particular expression has first been derived for the spherical scalar case in [19] and extended to the spherical vectorial case in [40, 36].
@@ 3.4.2 Arbitrary Phase Aberrations
More general aberration models can be introduced to describe imperfections in the optical system for PSF engineering or wavefront shaping. They can be introduced experimentally via a phase mask or a spatial light modulator at the pupil plane. These aberrations are often parametrized by Zernike polynomials (a set of orthonormal polynomials defined on the pupil disk) or a direct fixed phase mask to obtain desired PSFs. We write it in the most general case:
@@
Eq. (25) is composed of an inner product of the first $K$ Zernike polynomials and their corresponding coefficients $c_{k}$ and an additional term $W_{0}$ which can be used to include special cases not covered by the Zernike polynomials, e.g. a vortex phase ramp that leads to a donut PSF, typically used in stimulated emission depletion microscopy.
Note that arbitrary phase aberrations described in Eq. (25) may not necessarily be axisymmetric, hence, they can only be applied to the most general, Cartesian parametrization.
@@ 3.4.3 Apodization Factor
The apodization factor is an amplitude correction factor to ensure energy conservation during the change of basis from cylindrical coordinates (incident field $\mathbf{e}_{\mathrm{inc}}$) to spherical coordinates (far field $\mathbf{e}_{\infty}$), which matters especially for high-NA objectives. Since areas of cross-sections are modified, the field is also rescaled accordingly. Such rescaling ensures that the differential areas $dA_{1}$ on the plane and $dA_{2}$ on the sphere, as shown in Fig. 2, remain consistent under the change of coordinates. The corresponding correcting factor is
@@
when going from cylindrical to spherical in the focusing configuration of Fig. 1.
@@ 3.4.4 Gaussian Envelope
The incident illumination can also depart from a perfect uniform plane wave. In particular, we often assume a Gaussian envelope which can be expressed as
$A(\mathbf{s})=\exp\left\{-\frac{\sin^{2}\theta}{s_{\mathrm{env}}^{2}}\right\}$ (27)
@@ 4.1 Design and Usage
We provide a high-performance open-source PyTorch library psf-generator [^4] https://github.com/Biomedical-Imaging-Group/psf_generator to generate 2D and 3D PSFs. The library implements the four PSF models described in Section 3:
  • ScalarCartesianPropagator,
  • ScalarSphericalPropagator,
  • VectorialCartesianPropagator, and
  • VectorialSphericalPropagator.
In our unifying framework, these models encompass all previously proposed PSF models. Vectorial models are recommended for precise computations in high-NA scenarios. The Cartesian and spherical propagators perform equivalent integral computations, resulting in the same PSF. Both are proposed in our library as they differ in terms of computational efficiency and applicability. The spherical propagators are based on the axisymmetric assumption as described in Section 3.2. This includes all correction factors but the arbitrary phase aberrations introduced in Subsection 3.4.2.
@@
Users can choose between various propagator types, configure physical parameters (e.g., numerical aperture, wavelength, and field of view), numerical parameters (e.g., image dimensions and number of z-planes). Our library also allows the users to freely apply any kind of correction factors tailored for their microscope on any chosen propagator. The propagators use these parameters to first define the far field (pupil) and propagate it to obtain the focus field (PSF). Finally, the user can visualize, save and export the generated PSF using our built-in utility functions. Written in PyTorch, the library easily integrates into deep learning workflows, leveraging PyTorch’s native features such as automatic differentiation. We use the ZernikePy library [^5] https://pypi.org/project/ZernikePy/ to generate Zernike polynomials.
4.2 Code Optimization
@@
We have designed our library to follow the convention of PyTorch to benefit from its performance optimization to load data and parallelize computation. Namely, any tensor corresponding to a field has a shape of the form (z, channel, x, y), which corresponds to the typical PyTorch convention (batch, channel, x, y). Here, x and y correspond to the transverse sizes of 2D images. As computations are typically performed plane by plane, the batch dimension corresponds to the z axis. This enables efficient parallelization of Fourier transform for Cartesian models and numerical integration for spherical models. Finally, akin to grayscale versus RGB images, the channel dimension is equal to $1$ for scalar models and $3$ in the vectorial case.
@@
The Cartesian parametrization relies on multiple calls of 2D Fourier transforms with efficient sampling to obtain the desired field of view and pixel size. Padding zeros and using the Fast Fourier Transform (FFT) directly would be inefficient and significantly increase computational costs, as the physical sizes of pixels are typically very small in localization microscopy. We have implemented a custom variant of the 2D FFT in PyTorch, enabling arbitrary pixel size without additional computational cost. It is based on the chirp Z-transform which can be equivalently written as a convolution and computed using three FFTs. The computational complexity of a single plane is still $O(n\log(n))$ with $n$ the size of the transverse plane. The size of the Fourier transform is doubled as we only keep the valid convolution range. This method has been introduced in previous studies of the Cartesian models [16, 3, 29]. This approach also eliminates any artifacts caused by periodic boundary conditions, which are common with FFTs [3].
@@
On the other hand, the spherical parametrization relies on fast and accurate evaluation of multiple one-dimensional integrals that involve the Bessel function of first order $J_{0}$ over the interval $[0,\theta_{\max}]$ for a range of angles $\theta$ defined in Eq. (5) for the scalar case. To maximize speed, we vectorize the computation of the batch of 1D integrals at different defocus distances via torch.vmap. Then, for each 1D integral, we adopt the composite Simpson’s rule [43] to benefit from its high accuracy, good numerical stability and minimal computational overhead. This is not applicable to the Cartesian case because the disk-shaped support of the pupil function introduces a discontinuity, which prevents the application of numerical integration acceleration. More details and a comparison with other methods will be presented in Figure 6 and Section 5.2.2. The computational complexity in this case is merely $O(n)$ with $n$ the number of integration steps in the interval $[0,\theta_{\max}]$. Additionally, note that automatic differentiation of the Bessel functions is not natively supported by PyTorch as of version 2.3, so we also provide a differentiable version of the Bessel functions in our library as it is essential to learning-based tasks.

### v2

@@ 3.3 Vectorial Models
The electric field is a vectorial quantity. Consequently, vectorial propagation models are necessary to accurately account for the propagation and crosstalk between the components of the vector field. Precise vectorial models are crucial for high-NA systems, where the need to consider high angles arises.
In the vectorial model, the dependence on the incident field $\mathbf{e}_{\mathrm{inc}}(\mathbf{s})$ of the far field $\mathbf{e}_{\infty}(\mathbf{s})$ now requires us to perform the basis change from a cylindrical to a spherical coordinate system, as in
@@
$+$ $\left[\begin{array}[]{c}(-q_{s}+q_{p}\cos\theta)\sin 2\phi\\ q_{s}(1+\cos 2\phi)+q_{p}(1-\cos 2\phi)\cos\theta\\ -2q_{p}\sin\phi\sin\theta\end{array}\right]\frac{e_{\mathrm{inc}}^{y}(\theta,\phi)}{2},$ (13)
where $\mathbf{e}_{\mathrm{inc}}=[e_{\mathrm{inc}}^{x},e_{\mathrm{inc}}^{y},0]$. The so-called Fresnel transmission coefficients $q_{s}$ and $q_{p}$ have been introduced to account for the partial reflections at interfaces, which depend on the polarization state and incidence angle. For each polarization, the coefficients correspond to the product of all transmission coefficients for each interface from medium $m$ to $m+1$, with
$q_{s}^{m}=\frac{2n_{m}\cos{\theta_{m}}}{n_{m}\cos\theta_{m}+n_{m+1}\cos{\theta_{m+1}}}$
$q_{p}^{m}=\frac{2n_{m}\cos{\theta_{m}}}{n_{m+1}\cos\theta_{m}+n_{m}\cos{\theta_{m+1}}}.$ (14)
It is worth noting that in the low-NA limit ($\theta=0$) with $q_{s}=q_{p}=1$, (9) simplifies to $\mathbf{e}_{\infty}(\theta,\phi)=\mathbf{e}_{\mathrm{inc}}(\theta,\phi)$. This explains why the scalar model did not account for this geometric change of variables.
The Cartesian parameterization of the vectorial model consists of the integral
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{s_{x}^{2}+s_{y}^{2}\leq s_{\max}^{2}}\frac{\mathbf{e}_{\infty}(s_{x},s_{y})}{s_{z}}\exp{\left(\mathrm{i}ks_{z}z\right)}\exp{\left(\mathrm{i}k(s_{x}x+s_{y}y)\right)}\mathrm{d}s_{x}\mathrm{d}s_{y},$ (15)
which essentially boils down to the computation of the inverse Fourier transform of $(\mathbf{e}_{\infty}(s_{x},s_{y})\exp{\left(\mathrm{i}ks_{z}z\right)}/s_{z})$, similar to the scalar case.
With the help of coordinate transformations similar to the scalar case, we derive the spherical parameterization of the field in the focal region as
@@
$\exp{\left(\mathrm{i}k\rho\sin\theta\cos(\phi-\varphi)\right)}\exp{\left(\mathrm{i}kz\cos\theta\right)}\sin\theta.$ (16)
Inserting (9) into (3.3) and using the axisymmetric assumption of the incident field, we obtain a simplified expression of the focal field as
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2}\left[\begin{array}[]{c}[I_{0}^{x}-I_{2}^{x}\cos 2\varphi]-I_{2}^{y}\sin 2\varphi\\ -I_{2}^{x}\sin 2\varphi+[I_{0}^{y}+I_{2}^{y}\cos 2\varphi]\\ -2\mathrm{i}I_{1}^{x}\cos\varphi-2\mathrm{i}I_{1}^{y}\sin\varphi\end{array}\right],$ (17)
@@
$I_{0}^{a}(\rho,z)$ $=\int_{0}^{\theta_{\max}}e_{\mathrm{inc}}^{a}(\theta)\sin\theta(\cos\theta+1)\,J_{0}(k\rho\sin\theta)\exp{\left(\mathrm{i}kz\cos\theta\right)}\mathrm{d}\theta$
$I_{1}^{a}(\rho,z)$ $=\int_{0}^{\theta_{\max}}e_{\mathrm{inc}}^{a}(\theta)\sin^{2}\theta\,J_{1}(k\rho\sin\theta)\exp{\left(\mathrm{i}kz\cos\theta\right)}\mathrm{d}\theta$
$I_{2}^{a}(\rho,z)$ $=\int_{0}^{\theta_{\max}}e_{\mathrm{inc}}^{a}(\theta)\sin\theta(\cos\theta-1)\,J_{2}(k\rho\sin\theta)\exp{\left(\mathrm{i}kz\cos\theta\right)}\mathrm{d}\theta$ (18)
with $a\in\{x,y\}$.
@@ 3.4 Correction Factors
Precise PSF models commonly take into account several physical effects which may affect the true PSF. They are added as amplitude factors $a(\mathbf{s})$ and phase factors $W(\mathbf{s})$ in the original integral over solid angles in (1), leading to
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{\Omega}a(\mathbf{s})\exp{\left(\mathrm{i}W(\mathbf{s})\right)}\mathbf{e}_{\infty}(\mathbf{s})\exp{\left(\mathrm{i}k\mathbf{s}\cdot\boldsymbol{\rho}\right)}\mathrm{d}\Omega.$ (19)
Equation (19) enables us to express these correction factors in full generality for both the vectorial and scalar models and for both the Cartesian and spherical parameterizations. We present a detailed list of these correction factors in Section 3.4.1-3.4.4 and their graphical description in Figure 2.
Figure 2: Correction factors and their associated physical origins. Top part: phase correction factors, either introduced on the incident field or due to refraction. Bottom part: amplitude correction factors that model the incident beam envelope or the apodization factor for energy conservation.
3.4.1 Gibson-Lanni aberrations
Microscopes typically have stratified layers of different refractive indices. The biological sample is usually aqueous, on top of which we place a coverslip made of glass, and the whole sample is then put in a water or oil immersion medium to increase the numerical aperture. The microscope objectives are designed to provide aberration-free images in a specific setting with design values for refractive indices and thicknesses of each layer. Any mismatch introduces spherical aberrations due to refraction at the interface between the layers. These aberrations can be computed using the formula§§§Correction factors will be expressed in spherical coordinates, with $\sin\theta$ being computed in the Cartesian case as $\sin\theta=\frac{NA}{n_{i}}\sqrt{s_{x}^{2}+s_{y}^{2}}$.
@@ 3.4.1 Gibson-Lanni aberrations
$\quad\left.+t_{g}\sqrt{n_{g}^{2}-n_{i}^{2}\sin^{2}\theta}-t_{g}^{*}\sqrt{\left.n_{g}^{*}\right.^{2}-n_{i}^{2}\sin^{2}\theta}\right),$ (20)
where $n_{s}$, $n_{i}$, $n_{g}$ are the refractive indices of the sample, immersion medium, and glass, respectively, $t_{s}$, $t_{i}$, $t_{g}$ are the thicknesses of the sample, immersion medium, and glass respectively, and their counterparts with stars are the design conditions.
In practice, it is challenging to assess the thickness of the immersion medium. Since this distance is manually tuned to obtain an optimal focus of a point emitter at depth $t_{s}$ on the camera, this focusing condition gives the relation
@@
which can be inserted in 3.4.1. This particular expression has first been derived for the spherical scalar case in [19] and extended to the spherical vectorial case in [42, 36].
@@ 3.4.2 Arbitrary Phase Aberrations
Refined aberration models can be introduced to describe imperfections in the optical system, for the benefit of PSF engineering or wavefront shaping. Designed aberrations can be introduced purposefully via a phase mask or a spatial light modulator at the pupil plane. The aberrations are often parameterized by Zernike polynomials (a set of orthonormal polynomials defined on the pupil disk) or by a direct fixed phase mask to obtain desired PSFs. We write it in the most general case as
@@
Equation (22) is composed of the sum of two terms. The first term is an inner product of the first $K$ Zernike polynomials and their corresponding coefficients $c_{k}$. The second term can be used to include special cases not covered by the Zernike polynomials (e.g., a vortex phase ramp that leads to a donut PSF as is typically used in STED).
Since the arbitrary phase aberrations described in Equation (22) may not necessarily be axisymmetric, they can only be applied to the most general, Cartesian parameterization.
@@ 3.4.3 Apodization Factor
The apodization factor is an amplitude-correction factor that ensures energy conservation during the change of basis from cylindrical coordinates (incident field $\mathbf{e}_{\mathrm{inc}}$) to spherical coordinates (far field $\mathbf{e}_{\infty}$), which takes prominence especially for high-NA objectives. Since areas of cross-sections are modified, the field is also rescaled accordingly. Such rescaling ensures that the differential areas $dA_{1}$ on the plane and $dA_{2}$ on the sphere, as shown in Figure 2, remain consistent under the change of coordinates. The correcting factor is
@@
when going from cylindrical to spherical in the focusing configuration of Figure 1.
@@ 3.4.4 Gaussian Envelope
The incident illumination can also depart from a perfect uniform plane wave. In particular, we often assume a Gaussian envelope and expressed it as
$A(\mathbf{s})=\exp\left(-\frac{\sin^{2}\theta}{s_{\mathrm{env}}^{2}}\right),$ (24)
@@ 4.1 PyTorch library
We provide a high-performance open-source PyTorch library psf-generatorhttps://github.com/Biomedical-Imaging-Group/psf_generator to generate 2D and 3D PSFs. The library implements the four PSF models described in Section 3:
  • ScalarCartesianPropagator (3);
  • ScalarSphericalPropagator (5);
  • VectorialCartesianPropagator (15);
  • VectorialSphericalPropagator (17).
@@
Users can choose between these propagator types, as well as configure physical (e.g., numerical aperture, wavelength, and field of view), and numerical parameters (e.g., image dimensions and number of z-planes). Our library also allows the users to freely apply any kind of correction factors tailored to their microscope on any chosen propagator as described in Section 3.4.2. The propagators use these parameters to first define the far field (pupil) and propagate it to obtain the focus field (PSF). Finally, the user can visualize, save, and export the generated PSF using our built-in utility functions. Written in PyTorch, the library easily integrates into deep-learning workflows, leveraging native features of PyTorch, such as automatic differentiation.
In our unifying framework, these models encompass all previously proposed PSF models based on the Richards-Wolf integral. As discussed in Section 3.2, the spherical propagators require the axisymmetric assumption. When it is valid, the Cartesian and spherical propagators perform equivalent integral computations and result in the same PSF. Both are proposed in our library as they differ in terms of computational efficiency and applicability.
@@
Any tensor corresponding to a field has a shape of the form (z, channel, x, y) to benefit from PyTorch performance optimization, both for data loading and computation parallelization. Computation are performed plane-by-plane on the axial z axis, while x and y correspond to the transverse sizes of 2D images. Akin to grayscale versus RGB images, the channel dimension is equal to $1$ for scalar models and $3$ in the vectorial case. We use the ZernikePy libraryhttps://pypi.org/project/ZernikePy/ to generate Zernike polynomials.
@@
The Cartesian parameterization relies on multiple calls of 2D Fourier transforms with a desired field of view and pixel size. To account for the very small physical dimensions associated to pixels in localization microscopy, we have implemented a custom variant of the 2D FFT in PyTorch [16, 3, 29]. It enables arbitrary pixel sizes and is based on the chirp Z-transform, at the cost of a convolution and three FFTs. The computational complexity of a single plane is still $O(n\log n)$, with $n$ the size of the transverse plane. The size of the Fourier transform is doubled as we only keep the valid convolution range.
@@
The spherical parameterization relies on the fast and accurate evaluation of multiple one-dimensional integrals that involve the Bessel function of first order $J_{0}$ over the interval $0,\theta_{\max}]$, as defined in Eq. (5) for the scalar case. To maximize speed, we vectorize the computation of the batch of 1D integrals at different defocus distances via torch.vmap. Then, for each 1D integral, we adopt the composite Simpson’s rule [45] to benefit from its high accuracy, good numerical stability, and minimal computational overhead. The computational complexity in this case is $O(n)$, with $n$ the number of integration steps in the interval $0,\theta_{\max}]$. As automatic differentiation of Bessel functions is not natively supported by PyTorch as of version 2.3, we also provide a differentiable version of the Bessel functions in our library.
4.2 Open-source interoperability
Figure 3: Graphical user interface of our PSF simulation plugin integrated with napari (v0.4.0, 2025-06-17). Left: napari viewer panel enabling interactive 3D navigation and rendering options. Middle: interactive visualization area, displaying here a slice of a high-NA 3D PSF. Right: plugin interface for selecting the propagator model, specifying physical and numerical parameters (e.g., NA, wavelength, Zernike aberrations, CPU or GPU backend), and exporting results.
Motivated by FAIR principles for data and code management [46], we not only open-source our implementations but make them interoperable with other tools. Our goal is to further support the dissemination and adoption of accurate physical models for high-NA light focusing. We present two complementary examples: an intuitive and rapid visualization tool aimed at end users and an integration with a comprehensive optical simulation library intended for researchers.
napari is an open-source, Python-based image viewer designed for interactive visualization and analysis of multidimensional biological and scientific image data. Built for researchers and developers working with large imaging data from optical microscopy and other imaging modalities, it provides a high-performance interface for exploring, annotating, and processing images. napari is also extensible via plugins, benefiting from the scientific Python ecosystem.
We provide an interface to generate and visualize 3D PSFs using our PyTorch-based framework, as depicted in Figure 3. Available as a plugin on the napari hub******https://www.napari-hub.org/plugins/napari-psfgenerator, users can select between the four different propagation models and configure a set of physical and numerical parameters through a user-friendly graphical interface. This facilitates the introduction of high-NA correction factors to all four propagators types. PSFs are computed in real-time and rendered instantly, with support for both CPU and GPU backends. Our plugin allows immediate visual feedback and options to save the resulting volumes. It is designed to provide a code-free solution for the generation, visualization, and export of high-NA PSFs for downstream applications.
[…]
Chromatix is an open-source wave optics library capable of simulating a wide range of optical systems such as light field imaging, holography and quantitative phase imaging [38]. Built on JAX, it supports GPU acceleration and full end-to-end differentiability, making it particularly suitable for large-scale inverse problems and integration with deep learning applications. Compared to PyTorch, JAX offers more seamless support for automatic differentiation and just-in-time (JIT) compilation, which can be advantageous for optimizing physics-based simulations based on wave optics.
To ensure maximal performance and seamless integration with the rest of the library, the high-NA focusing models have been reimplemented in JAX. We focused on the vectorial Cartesian parametrization, which offers greater generality than the spherical formulation and supports arbitrary pupil fields and aberrations. This functionality is introduced via the high_na_ff_lens() function, which complements the existing ff_lens() implementation used to model the Fourier transform of low-NA lenses. This addition enables Chromatix to accurately model high-NA microscope systems and facilitates integration into existing simulation pipelines.
These two examples highlight the importance of actively promoting interoperability within the open-source ecosystem. Nevertheless, all figures and benchmarks presented in this work were produced based on our comprehensive PyTorch-based library, which serves as the foundation of our study.